OpenStudy (anonymous):

f"(x)=4x-3 and f'(-3)=-1 and f(-3)=0 find f'(x) and f(1)

OpenStudy (anonymous):

do you know integrals?

OpenStudy (anonymous):

is that the same as antiderivatives

OpenStudy (anonymous):

f'(x)=(4(x^2))/2 + 3x + c= 2x^2 + 3x + c then use f'(3)=-1=2(3)^2 + 3(3)+c to solve for c=-28 f(x)=(2/3)x^3 + (3/2)x^2 +28x +c

OpenStudy (anonymous):

now see if you can use f(-3)=0 to find c again and then find f(1)

OpenStudy (anonymous):

i got the same answer but it tells me that f'(x) is wrong...